The uniform disc has a MOI of 0.6kg-msq around the axis that passes through the centre and is perpendicular to the plane of the page.If a segment is cut from the disc, making an angle of 60deg with the centre of disc, what is the MOI of the remaining part?


DC Pandey(pg--9 part 2--Rotational Motion)

Grade:11

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Shirsa Udgata,

Ans:-See the orientation is remained unchanged. Hence the final MI is also =1/2 M′r² but here M′ is the mass of the remaining part=5/6(0.6)=0.5 Kg-m²

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Shirsa Utgata !!!

 


Regards,

Askiitians Experts

Soumyajit IIt-Kharagpur